OpenStudy (anonymous):

(x-1/2)(4x^3-2x^2+5x-6) this is what i got, does it look right? 4x^4-8/2x^3+12/2x^2-17/2x+6/2

hartnn (hartnn):

you can simplify that

hartnn (hartnn):

8/4 =... ? 12/2=.. ? 6/2 =... ?

hartnn (hartnn):

everything else is correct :)

OpenStudy (anonymous):

i thought about that but i was not sure if i should or not

hartnn (hartnn):

yes, since you can, you should :)

OpenStudy (anonymous):

got it, thanks

hartnn (hartnn):

welcome ^_^